# Advancing Mobile GUI Agents: A Verifier-Driven Approach to Practical Deployment

### v2

@@ 6.3. Training Scaling Law in V-Droid
V-Droid improves with more training data. Fig.11(a) and Fig.11(b) illustrate how the V-Droid’s performance scales with increasing training data. On AndroidWorld, the success rate improves from 15.0% to 37.5%, then further scales to 47.4% and 59.5% as the number of training data pairs increases from 9K (Human-Annotated) to 110K (Human-Agent Joint-Annotated). Similarly, on the MobileAgentBench, the success rate increases from 17.0% to 49.0% as more training data becomes available. This continuous improvement highlights that V-Droid benefits from diverse app environments, instructions, and execution trajectories, which gradually expand throughout the training process. The increasing dataset variety enhances the agent’s generalization ability and robustness, leading to more effective decision-making across different tasks.
Human annotation overhead decreases with better verifier. As shown in Fig. 12(a), the human annotation effort, measured as the ratio of data pairs collected by human annotators, gradually decreases across multiple iterative annotation and training cycles. After training with 27K and 55K data pairs, the AUC improves from 0.55 to approximately 0.8, demonstrating a significant enhancement in the verifier’s ability to predict decision correctness. This improved accuracy enables human annotators to quickly rectify errors and efficiently recover the data collection process with minimal overhead. To further enhance the verifier’s capabilities, we continue to scale the dataset using the human-agent joint annotation scheme in V-Droid, ensuring progressive improvements in annotation efficiency and model performance.
6.4. Comparison with Design Alternatives
Architecture and Training Alternatives. In addition to the comparison with baselines, we also craft three baselines to justify the designs in V-Droid, including LLM-as-a-Judge, Selector and Generator. LLM-as-a-Judge follows the architecture of V-Droid, but the verifier is replaced with GPT-4. The action score is obtained by extracting the logits value of ”Yes” from the output. Selector is presented with the HTML descriptions along with the extracted action lists and is fine-tuned to output the correct action number. Generator follows the architectures of T3A (Rawles et al., 2024) but replaces the policy agent with a fine-tuned LLama-3.1-8B that is trained to output thought chains followed by the chosen actions. We adopt supervised fine-tuning for both design. The number of tasks completion trajectories used for training and the training setting is kept the same. The experiment is conducted with on $55$K training samples.
@@ 6.4. Comparison with Design Alternatives
Table 2. Comparing V-Droid with the design alternatives of agent architectures and corresponding training approaches with training data from $1.1$K task steps.
@@
Table 2 highlights the limitations of LLM-As-A-Judge using Llama-3.1-8B without training, which fails to assign accurate action scores, resulting in a 0% success rate. While GPT-4 performs better due to stronger reasoning and instruction-following abilities, it also outperforms its own generator-based approach, reinforcing the generation-verification gap. However, it is observed that GPT-4 often assigns high scores to multiple actions or low scores to all actions, exposing a misalignment between token space and action space. Training Llama-3.1-8B as either a selector or generator improves success rates to 35.8% and 27.4%, respectively. However, these models still underperform compared with V-Droid, which leverages a verifier-driven workflow with $P^{3}$ training. The advantage of $P^{3}$ training comes from its pair-wise learning structure, where each step with $N$ available actions generates $N-1$ training pairs, effectively amplifying the training data by $N-1$ times. Additionally, pair-wise training enhances the model’s ability to distinguish similar UI elements and actions, a crucial factor for accurate action selection.
In Table 2, LLM-as-a-Judge based on Llama-3.1-8B without training cannot assign accurate scores for action selections, which leads to 0% success rate. The performance of LLM-as-a-Judge of GPT-4 is not only better compared with Llama-3.1-8B due to better reasoning and instruction following, but also outperforms than using it as a generator, which validates the generation-verification gap (Lightman et al., 2023; Zhang et al., 2024a). However, we observe that it assigns high scores to multiple actions or low scores to all actions, which reveals the gap between token space and the action space. While training Llama-3.1-8B as a selector or generator improves their success rate to 35.8% and 27.4%, their performance are worse compared with V-Droid with the verifier-driven workflow using $P^{3}$ training. The reason would be that $P^{3}$ training provides $N-1$ pairs for one step if there are $N$ available actions, which augments the step-wise data by $N-1$ times. The pair-wise training also allows the model to learn to distinguish similar UI and actions, which are important for accurate action selection.
@@
Table 3. Decision Alternatives on Constructing Working Memory.
Design Alter SR Latency (s)
LLM-based Memory 59.5% 3.03
Rule-based Memory 46.1% 1e-5
Action History Only. 40.0%
@@ 6.5. System Overhead
Running the verifier of V-Droid on GPUs requires approximately $16$GB of memory, including the KV cache. Across the evaluated benchmarks, V-Droid verifies an average of $50.3$ actions per task step.
Such verifications could be significantly accelerated with the batching inference and prefix caching. We highlight the decision-making latency on 4×4$\times$4 × NVIDIA GTX A100 80G, 4×4$\times$4 × NVIDIA GTX A6000, and 2×2$\times$2 × NVIDIA GTX 4090.
As shown in Table 4, given the optimization of the prompt format in V-Droid that maximizes the length of the shared prefix, a 10×10$\times$10 × speed up is obtained from the prefix caching across different actions, steps, and tasks. Furthermore, V-Droid turns the auto-regressive decision making scheme of LLM agents into the parallel pre-filling only verification scheme, which can be conducted in batch on multiple GPUs. The parallelism further decreases the selection time to 0.42s, 0.44s, and 0.52s with 4×4$\times$4 × A100, 2×2$\times$2 × 4090, and 4×4$\times$4 × A6000, respectively.
We also try to infer the overhead running the verifier on mobile devices. Taking the shared prefix into account, the total input token count for all action verifications per step is approximately 1.1K tokens. Given a prefill speed of 450 tokens per second on the Qualcomm Snapdragon 8 Gen 3 NPU (Project, 2024), the decision-making latency could be as low as 2.5 seconds.
Table 4. Decision-making latency with and without the prefix caching (PC) on different instances of 4090, A100 and A6000 GPUs.
@@ 7. Discussion
Test-time scaling. Recent advancements in slow-thinking techniques have demonstrated their potential to enhance the reasoning capabilities of LLMs (Guan et al., 2025). These techniques are fully compatible with V-Droid; for instance, additional thinking tokens could be generated before assigning scores, which would be particularly beneficial for critical steps where V-Droid exhibits uncertainty, as indicated by score entropy (detailed in §5.1). We plan to explore this direction in future work.
@@ 8. Conclusion
For the first time, V-Droid demonstrates near-real-time, effective decision-making for mobile agents. It discretizes the decision space for mobile interactions into a finite set of action candidates, transforming the autoregressive process of generator-based agents into a parallelized, prefilling-only scoring mechanism using a generative verifier. $P^{3}$ training, incorporating self-correction pairs and a scalable human-agent joint annotation framework, significantly enhances the verifier’s decision-making capabilities. Experimental results showcase the training scaling law of the verifier-driven architecture for mobile task automation.

### v3

@@ 6.3. Training Scaling Law in V-Droid
V-Droid improves with more training data. Fig.12(a) and Fig.12(b) illustrate how the V-Droid’s performance scales with increasing training data. On AndroidWorld, the SR improves from 15.0% to 37.5%, then further scales to 47.4% and 59.5% as the number of training data pairs increases from 9K (Human-Annotated) to 110K (Human-Agent Joint-Annotated). Similarly, on the MobileAgentBench, the SR increases from 17.0% to 49.0% as more training data becomes available. This continuous improvement highlights that V-Droid benefits from diverse app environments, instructions, and execution trajectories, which gradually expand throughout the training. The increasing dataset variety enhances the agent’s generalization ability and robustness, leading to more effective decision-making across different tasks.
Human annotation overhead decreases with better verifier. As shown in Fig. 13(a), the human annotation effort, measured as the ratio of data pairs collected by human annotators, gradually decreases across multiple iterative annotation and training cycles. After training with 27K and 55K data pairs, the AUC improves from 0.55 to approximately 0.8, demonstrating a significant enhancement in the verifier’s ability to predict decision correctness. This improved accuracy enables human annotators to quickly rectify errors and recover the data collection process with minimal overhead. To further enhance the verifier’s capabilities, we continue to scale the dataset using the human-agent joint annotation scheme in V-Droid, ensuring progressive improvements in annotation efficiency and model performance.
@@
6.4. Comparison with Design Alternatives
Architecture and Training Alternatives. We craft three baselines to justify the designs in V-Droid, including LLM-as-a-Judge, Selector and Generator. LLM-as-a-Judge follows the architecture of V-Droid, but uses GPT-4 as the verifier. The action score is obtained by extracting the output logits value of ”Yes”. Selector is presented with XML descriptions and the extracted action lists and is fine-tuned to output the correct action number. Generator follows the architectures of T3A (Rawles et al., 2024) but replaces the policy agent with a fine-tuned LLama-3.1-8B that is trained to output thought chains followed by the chosen actions. We adopt supervised fine-tuning for both designs.
Table 3. Comparing V-Droid with the design alternatives of agent architectures and corresponding training approaches with training data from 1.11.1K task steps.
@@ 6.4. Comparison with Design Alternatives
Table 3 highlights the limitations of LLM-As-A-Judge using Llama-3.1-8B without training, which fails to assign accurate action scores, resulting in a 0% success rate. While GPT-4 performs better due to stronger reasoning and instruction-following abilities, it also outperforms its own generator-based approach, reinforcing the generation-verification gap. However, it is observed that GPT-4 often assigns high scores to multiple actions or low scores to all actions, exposing a misalignment between token space and action space. Training Llama-3.1-8B as either a selector or generator improves success rates to 35.8% and 27.4%, respectively. However, these models still underperform compared with V-Droid, which leverages a verifier-driven workflow with $P^{3}$ training. The advantage of $P^{3}$ training comes from its pair-wise learning structure, where each step with $N$ available actions generates N1N-1 training pairs, effectively amplifying the training data by N1N-1 times. Additionally, pair-wise training enhances the model’s ability to distinguish similar UI elements and actions, a crucial factor for accurate action selection.
@@
Table 4. Decision Alternatives on Working Memory.
Design Alter. SR Latency (s)
LLM-based 59.5% 3.03
Rule-based 46.1% 1e-5
Actions History 40.0%
@@ 6.5. System Overhead
Running the verifier of V-Droid on GPUs requires approximately 1616GB of memory, including the KV cache. Across the evaluated benchmarks, V-Droid verifies an average of $50.3$ actions per task step.
Such verifications could be significantly accelerated with the batching inference and prefix caching. We highlight the decision-making latency on $4\times$ NVIDIA GTX A100 80G, $4\times$ NVIDIA GTX A6000, and $2\times$ NVIDIA GTX 4090.
As shown in Table 5, given the optimization of the prompt format in V-Droid that maximizes the length of the shared prefix, a $10\times$ speed up is obtained from the prefix caching across different actions, steps, and tasks. Furthermore, V-Droid turns the auto-regressive decision making scheme of LLM agents into the parallel pre-filling only verification scheme, which can be conducted in batch on multiple GPUs. The parallelism further decreases the selection time to 0.42s, 0.44s, and 0.52s with $4\times$ A100, $2\times$ 4090, and $4\times$ A6000.
We also try to infer the overhead running the verifier on mobile devices. Taking the shared prefix into account, the total input token count for all action verifications per step is approximately 1.1K tokens. Given a prefill speed of 450 tokens per second on the Qualcomm Snapdragon 8 Gen 3 NPU (Project, 2024), the decision-making latency is around 2.5 seconds.
Table 5. Decision-making latency with and without the prefix caching on different instances of 4090, A100 and A6000 GPUs.
@@
6.6. Failure Study
We conducted a manual analysis of the failure cases for V-Droid and MobileGPT across more than 300 tasks spanning three benchmarks. We identify four categories of failure: 1) Hallucinated decisions, where the agent selects an incorrect action despite being provided with the correct context; 2) Inaccurate memory, where erroneous contextual memory leads the agent to make wrong decisions; 3) Incomplete actions, wherein available action types are insufficient to complete the task; and 4) Modality limitations, where tasks cannot be accomplished by agents based solely on text-only LLMs.
As illustrated in Fig. 14, failure cases of MobileGPT primarily stem from suboptimal decision-making. For example, given the task instruction Add one expense (\$307.01, in Health Care categories) to APP Expense Pro, MobileGPT erroneously inputs the wrong amount and neglects to select the Health Care category, whereas V-Droid accurately provides all required information. This observation further underscores the necessity of post-training for GUI agents for a larger scale. Additional sources of failure include an incomplete action space caused by inaccurate information from the accessibility service, the lack of visual processing required for interpreting images and videos.
@@ 8. Conclusion
For the first time, V-Droid demonstrates near-real-time, effective decision-making for mobile agents. It discretizes the decision space for mobile interactions into a finite set of action candidates, transforming the autoregressive process of generator-based agents into a parallelized, prefilling-only scoring mechanism using a generative verifier. $P^{3}$ training and the scalable human-agent joint annotation framework, significantly enhances the verifier’s decision-making capabilities. Experimental results showcase the training scaling law of the verifier-driven architecture for mobile task automation.